A groundwater contains a dissolved oxygen concentration of 5 mg/L. How many liters of groundwater would it take to oxidize 1 g of toluene to CO2?
Answer #1

if we assume the break of the ring

\\ C_7H_8+9O_2\rightarrow 7CO_2 + 4H_2O\\ 92g.........288g\\ 1g............x

x=\frac{1*288}{92}=3.13g

\\ 5*10^{-3}g...........1L\\ 3.13g...............V

V=\frac{3.13g*1L}{5*10^{-3}g}=626L